OpenStudy (anonymous):

3x+12y=213 9x+6y=189 these are system story problems. im stuck on the "show your work"

OpenStudy (anonymous):

i need some help here!!!

OpenStudy (anonymous):

will anyone help me?

OpenStudy (arbershabani97):

you multiply second equation with 2, and by that you get 12y, that makes the Y's cancel out, and you solve with substitution

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now what

OpenStudy (arbershabani97):

I'll just start answering the equation and then you can continue it, just a minute

OpenStudy (arbershabani97):

3x+12y=213 (9x+6y=189)*2 3x+12y=213 -(18x+12y=378) -15x=-165 and then you can do it by yourself, if you don't understand anything, ask freely

OpenStudy (anonymous):

ok thankyou for your help

OpenStudy (anonymous):

have a great day

OpenStudy (arbershabani97):

you too! :)